# CAVE: Controllable Authorship Verification Explanations

### v1

@@ 3.4 Empirical Results
Table 3 shows the empirical comparison between the baselines and Cave. We compare all the results in terms of Accuracy and analyze Cave’s results for Consistency.
@@
We first look at the LLaMa-3-8B models trained on individual datasets. For IMDb62 and FanFiction we note that Cave beats all baselines (including GPT-4-Turbo which is orders of magnitude larger) on both accuracy and consistency. For Blog-Auth we note that Cave beats 3 out of the 6 baselines in terms of accuracy, and is competitive with the remaining baselines (60.4% as compared to 60.7%, 62.0%, 62.2%). Blog-Auth’s model also beats all relevant baselines in terms of consistency.
@@
Inspired by the above results, we decided to train a model on the set of three datasets instead of a single dataset to get a better performance across the board. We train LLaMa-3-8B with IMDb62, Blog-Auth and FanFiction training sets together (the last row in Table 3). We note that this model gives the best consistency for all datasets. Further, it beats all baselines for IMDb62 and FanFiction in terms of accuracy, and offers a competitive accuracy for Blog-Auth. There is minimal loss in accuracy when we move from single dataset to multi-dataset training, which is complemented by increased consistency.
@@
special characters style, capitalization style 35 37 32
@@
Table 4: Human Pilot: This table shows the number of examples out of 60 where all 3 annotators agree that the rationale conforms to the property being analyzed. P1 - ‘Detail-Consistency’, P2 - ‘Factual-Correctness’, P3 - ‘Label-Consistency’. We report these numbers for all features except punctuation (discussed in Section 4.3).
3.5 Human Pilot
Prior works in the field of Av largely focus on task accuracy, with much lesser focus on explainability Patel et al. (2023) and by extension, human evaluation of the explanations. In this work, we take the first step towards human analysis of Av explanations by providing a pilot study on our model Cave’s rationales. We pick $20$ document pairs per dataset (total of $60$ pairs for analysis) and evaluate the rationales generated by the multi-dataset Cave (last row in table 3). We use this study to understand the utility of these rationales to downstream applications Joshi et al. (2023) and to learn about potential areas for improvement van der Lee et al. (2019). We analyze the generated rationales (for each linguistic feature as in Table 2) via three axes as described below; further, given the high complexity of this task, we opt to do an in-house human evaluation, with three distinct annotators per dataset. Note that during human evaluation, we also collect information about what is wrong with the rationale if an annotator marks the property as unsatisfied.
  1. 1.
    Detail-Consistency: Are the details mentioned in the rationale consistent with the input documents? Or are they hallucinated? (eg:- if the rationale mentions that both documents use parantheses, but the documents don’t, then the details are hallucinated)
  2. 2.
    Factual-Correctness: Are the rationales factually correct? (eg:- if the rationale says that the text was informal in tone, while it was actually formal, or if it misinterprets an acronym to be author-specific slang, etc. it is incorrect)
  3. 3.
    Label-Consistency Cave’s output format has a distinct rationale and intermediate label (YES/NO/MAYBE) for each linguistic feature. As defined in Section 2.2, we use these intermediate labels to check the overall consistency of the rationale with the final label. In this human evaluation however, we check whether each individual rationale is consistent with its corresponding label.
In Table 4, we show number of examples (out of 60) for which the three annotators agree that the rationale aligns with the property being analyzed. We report numbers for all linguistic features except for punctuation style, since the FanFiction dataset suffers from a dataset bias error for punctuation (discussed in Section 4.3). Out of the 21 pairs of linguistic features (7) x properties (3) analyzed in Table 4, annotators find that for 15 of them, $\geq 70\%$ of the datapoints are aligned with the property analyzed. For the cases where human annotation finds the rationales to be lacking (such as ‘capitalization’, ‘acronym’ analysis for all datasets) we take a deeper look into the samples and the annotators’ comments. We find that errors are split between data-bias type errors (such as for FanFiction punctuation, where the dataset uses double instead of single quotes for apostrophes, leading to persistent errors across all examples) and genuine errors on the part of the model (such as hallucinating punctuation marks that weren’t present in the documents). We also note that there were a number of cases where annotators marked that ‘YES’ or ‘NO’ labels (intermediate) should be marked as ‘MAYBE’ and vice versa - we discuss the nuances of this through examples in Section 4.3.
We report these numbers for individual datasets in Table 13 in Appendix D. We also release the complete set of test-set rationales generated by Cave 777docs.google.com/spreadsheets/d/1npLyOAwdnIAWjmQbCip5w_Ajv34YU1v7uiUhp4CgmPc/edit?usp=sharing.
3.6 Out-of-domain inference
Test $\rightarrow$ IMDb62 Blog-Auth FanFiction
Train $\downarrow$ Acc. Cons. Acc. Cons. Acc. Cons.
IMDb62 57.7 0.95 58.7 0.92
Blog-Auth 73.8 0.93 50.6 0.74
FanFiction 68.6 0.70 57.4 0.83
Table 5: Out of domain performance by single-dataset Cave models
4 Discussion
4.1 Accuracy vs. Explainability
Early works on self-rationalization such as Kojima et al. (2022); Marasovic et al. (2022); Wei et al. (2022) showed that the generation of rationales leads to improved task accuracies; however, most such works place a higher focus on task performance than on the quality of the rationale. Recent works such as Joshi et al. (2023); Ramnath et al. (2023); Lyu et al. (2023) discuss the need for increased quality of the rationale in terms of properties such as faithfulness, plausibility, etc. in order for them to be truly useful and trusted in downstream applications. Hence, in our efforts towards better authorship verification, we place a high importance on the explainability and faithfulness of rationales; by fine-tuning a smaller model with training data that displays both consistency and task accuracy, we aim to find a balance between explainability and accuracy. The zero-shot application of our method (i.e., Cave’s prompt with GPT-4-Turbo) improves in consistency but is lesser in task accuracy as compared to prior works. We achieve our goals of balancing both metrics by filtering data before distilling it. Our final model Cave leads to both increased explainability (via consistency of labels) as well as increased/competitive task accuracies. In addition, the availability of intermediate labels allows for consistency checking at run time—that is, we can automatically identify to an end user whether or not an explanation is likely to be reliable, providing them with another important tool as they are analyzing potentially complex authorship problems.
4.2 Efficiency & Security
While GPT-4-Turbo performs reasonably well with Cave’s prompt, we desire to have a small, on-server model to achieve our goals of lower computational/financial cost and more security for users’ sensitive data. At this stage, we had two options regarding smaller LMs: finetune a pretrained LM, or directly use an instruction-finetuned LM.
We see in Table 3 that as we move from GPT-4-Turbo to LLaMa-3-8B-Instruct (which is a significant drop in model size), the accuracy drop faced by Cave’s prompt is much higher than that faced by CoT and PromptAV. This necessitates us to train a LLaMa-3-8B for this task.
@@ 4.2 Efficiency & Security
4.3 We measured consistency. What’s next?
@@
In this work, we proposed Cave that is trained to have consistency between the different steps of the rationale and the label. However, in addition to this, there are a number of properties that are necessary for the rationale to be of good quality Joshi et al. (2023); Ramnath et al. (2023). Some of these are properties that are easily verifiable by humans, even if there exist no prevalent automatic methods to measure them: for example, factual correctness of the rationale, hallucinated details in the rationale, etc. that we measured in our human evaluation (Section 3.5). There are also properties that are hard to verify, even by human evaluators: for example, completeness of the rationale with respect to its analysis (i.e., has the rationale missed any significant similarity/dissimilarity between the documents?). In fact, properties like completeness do not have a comprehensive implementation even for simpler self-rationalization tasks like in multi-choice QA Ramnath et al. (2023). In the future, we would like to work on reward modeling for these properties for self-rationalization in Av; we hope to develop widely usable metrics for these properties.
In line with this discussion, we extend our human evaluation from Section 3.5 to provide qualitative examples. Our highest number of errors (ascertained from human evaluation) were hallucinated details in the rationale, particularly in the case of punctuation. We found that for a number of examples, our model cited the presence of punctuation marks such as ellipses, parantheses, etc. as the similarity/difference between documents, when in reality, those did not exist in the document. We report commonly observed errors in Table 6. Dataset errors (such as double quotes being used as apostrophes in FanFiction) affect the judgement of the model. Further, the model sometimes uses topic of the text as a reason to indicate or refute authorship between two documents; topic cannot be an absolute factor in deciding authorship Wegmann et al. (2022a), and hence such labels must be ‘MAYBE’ - this was the second highest number of errors as noted by annotators. As future work, we would like to address these shortcomings with targeted training. We would also like to look at dynamic weighing of linguistic features at inference time, i.e., which features contributes more/less to the authorship analysis of these two documents? Finally, we also discuss random errors owing to computational limitations (such as limited context window in language models) in Appendix E.
@@ 4.4 Potential Risks
As we discussed in Section 1, the task of authorship verification is high-stakes in nature. While we strive to enforce explainability in our model Cave, it is still a language model which is trained on data generated from the black-box LM GPT-4-Turbo; both GPT-4-Turbo’s and LLaMa-3-8B’s own biases could have been retained in Cave. Hence, we caution downstream users to keep this in mind; Cave’s outputs should be used only after due understanding and verification. We discuss more details in the Limitations section at the end.
@@ 5 Conclusion
Existing methods for authorship verification face the lack of accessible explainability. In this work, we train an in-house language model Cave to generates structured, consistent explanations for Av. We present strong experimental results on three difficult Av datasets. We also present a human pilot study on the quality of generated rationales. Finally, we provide detailed discussions on explainability in Av and explore errors made by Cave due to factors such as hallucination, dataset biases, etc. As future work, we would like to explore reward metrics and reward-based learning for Av explanations.
@@ Acknowledgements
Like any large language model, Cave can unintentionally lead to toxic, incorrect or hallucinated text; it is up to user to use the model mindfully and responsibly. Further, since Cave is designed for the task of Av which is extremely sensitive in nature, it is the responsibility of the user to duly verify the data used to train the model, and also verify the output given by the model if/before using it. Further, as noted in the main paper, our models have not been explicitly trained for completeness of rationale; that is, apart from wrong information the generated explanation may have, it could also have missed some crucial similarity/dissimilarity between the documents.
Further, the Cave models trained in this paper were trained on data sampled from yet another language model GPT-4-Turbo and not on human-written gold standard data; it is possible that harmful as well as harmless biases that GPT-4-Turbo (by process of its training and its training data) might have been distilled into our model. It is necessary to keep this in mind while using our training data or our models.
Lastly we note that our method as such has practical limitations faced by any NLP task; our models are restricted by the size of the context window allowed by model (both input and output text) (also keeping in mind restrictions due to computational costs, financial costs and GPU size available). This can lead to data/reasoning missed by the model in its reasoning.
@@ 8. Other Aspects:
PromptAV (and by extension, CoT) which serve as the inspiration for Cave produce outputs which cannot be automatically checked for consistency; these prompts produces a free-text explanation followed by a final score which can be thresholded to produce a YES/NO Av label. As observed in the experiments of Hung et al. (2023), models tend to produce low confidence scores with PromptAV, which means that for better accuracy, the optimal threshold in their experiments with GPT-3.5-Turbo was found to be 0.2-0.3, which will most probably not be consistent with the free text rationale (that is, if the rationale says that it thinks it is not confident that the authors are the same, and the confidence is 0.25, the automatic label prediction will still assume that the predicted label is YES). Further, with just the score as the label, it is still possible that we misconstrue the label because the confidence score and the model’s predicted label in the text are different, as shown in the examples in Tables 9 and 10.
@@ Appendix C Hyperparameters
Appendix D Human Evaluation

### v2

@@ 3.4 Empirical Results
Table 3 shows the empirical comparison between the baselines and Cave. We report Accuracy for all models, and Consistency for Cave.
@@
We first look at the LLaMa-3-8B models trained on individual datasets. For IMDb62, Cave beats all baselines (including zero and few-shot GPT-4-Turbo, which is orders of magnitude larger in parameter size) on both accuracy and consistency; for FanFiction, Cave beats all baselines except for few-shot GPT-4-Turbo. For Blog-Auth, Cave beats significant baselines such as zero-shot GPT-4-Turbo, and is competitive with the remaining baselines in terms of both metrics.
@@
We also train a multi-dataset Cave on all 3 datasets together to observe any transfer-learning trends (the last row in Table 3). This model gives the best consistency for all datasets. There is minimal loss in accuracy when we move from single dataset to multi-dataset training, which is complemented by increased consistency; for IMDb62, multi-dataset Cave still beats all baselines in terms of both metrics.
Why show two zero/few-shot baselines?
In addition to the strong (online) baseline of GPT-4-Turbo, we also show zero/few-shot results using a less-strong local model, LLaMa-3-8B-Instruct. The significant performance degradation there (for all cases except few-shot Blog-Auth) shows that it is insufficient merely to take existing approaches and run them using a local instruction-tuned model; our proposed distillation strategies are essential to get a high performance with local, offline models that face the disadvantage of having a much lower parameter size.
LLaMa-3-8B for Cave.
In Appendix D, we report results on Cave with other offline student models Jiang et al. (2023); Team (2024); Stability-AI (2024) of varying sizes to inform design choices in downstream applications; we find that LLaMa-3-8B is the best performing Cave, with Mistral-7B Jiang et al. (2023) being the second-best.
3.5 Human Pilot
Prior works in this field largely focus on task accuracy, with much lesser focus on explainability Patel et al. (2023) and by extension, human study of the explanations. In this work, we take the first step towards human analysis of Av explanations by providing a pilot study on Cave’s rationales.
In-house human study: Av + explanations is a highly complex task; annotators require a considerable amount of training and practice to be skillful in their evaluation. Hence, for our evaluation, we opted to conduct a small but high-quality in-house human study, instead of a large-scale study with platforms such as Amazon MTurk.
We pick $20$ document pairs per dataset (total of $60$ pairs) and evaluate the rationales generated by multi-dataset Cave. We analyze the strengths & weaknesses of Cave’s rationales, which help to understand the utility of these rationales to downstream applications Joshi et al. (2023) and to learn about potential areas for improvement van der Lee et al. (2019). We analyze the generated rationales (for each linguistic feature as in Table 2) via three properties as described below; we employ 3 distinct annotators per datapoint. We also require the annotators to comment on why a rationale is unsatisfactory, if they find it to be so with respect to any property; this helps to perform a more detailed analysis at the instance level.
1. Detail-Consistency: Are the details mentioned in the rationale consistent with the input documents, or are they hallucinated? (eg:- if the rationale mentions that both documents use parantheses, but the documents don’t, then the details are hallucinated)
2. Factual-Correctness: Are the rationales factually correct? (eg:- if the rationale says that the text was informal in tone, while it was actually formal, or if it misinterprets an acronym to be author-specific slang, it is incorrect)
3. Label-Consistency Cave’s output format has a distinct rationale and intermediate label (YES/NO/MAYBE) for each linguistic feature. As defined in Section 2.2, we use these intermediate labels to check the overall consistency of the rationale with the final label. In this human evaluation however, we check whether each individual rationale is consistent with its corresponding label.
@@ 3.5 Human Pilot
special characters & capitalization style 35 37 32
@@
Table 4: Human Pilot: This table shows the number of examples out of 60 where all 3 annotators agree that the rationale conforms to the property being analyzed. P1 - ‘Detail-Consistency’, P2 - ‘Factual-Correctness’, P3 - ‘Label-Consistency’. We report these numbers for all features except punctuation (discussed in Section 3.5.
@@
In Table 4, we show number of examples (out of 60) for which the three annotators agree that the rationale aligns with the property being analyzed. We also show aggregate statistics for each property across all linguistic features, i.e., $\frac{sum(P1)}{60*7}$ where 7 is the number of linguistic features being analyzed. We report numbers for all linguistic features except for punctuation style, since the FanFiction dataset suffers from a dataset bias error for the same (discussed further below).
Aggregate Analysis. Annotators find that Cave’s rationales align with properties P1 and P2 in 81.4% of the cases, and with property P3 in 70.4% of the cases. That is, Cave on average hallucinates or presents incorrect details in $\sim$19of the cases, and produces inconsistent intermediate labels in $\sim$30of the cases.
Fine-grained analysis. Out of the 21 pairs of linguistic features (7) x properties (3) analyzed, annotators find that $\geq 70\%$ of the datapoints are aligned with the property analyzed for 15 pairs. For the cases where annotators finds the rationales to be dissatisfactory (such as ‘capitalization’, ‘acronym’) we take a deeper look into annotators’ comments.
Errors due to dataset biases. Annotators noticed that documents in the FanFiction dataset had a bias - they all used double instead of single quotes for apostrophes (eg:- it“s instead of it‘s), leading to persistent errors across all examples - Cave concluded that these double quotes were indicative of authorship, and consistently marked that FanFiction document pairs had the same author with respect to punctuation style.
Qualitative examples & Annotator comments. We report commonly observed rationale errors in Table 5. Hallucination of details was a common error, especially for punctuation style and special characters / capitalization; for example, Cave cited the presence of punctuation marks such as ellipses, parantheses, etc. as the similarity/difference between documents, when in reality, those did not exist in the document. One hypothesis is that LLMs are bad at these linguistic features since they deal with individual characters and not broad pieces of text like the other features. Another common error was the Cave’s usage of topic of texts as a reason to indicate or refute shared authorship of two documents; topic cannot be an absolute factor in deciding authorship Wegmann et al. (2022a), and hence such labels must be ‘MAYBE’. There were also cases where Cave used trivial grounds to reason about authorship (for example, ‘documents have the same author because both have standard acronyms and capitalization’, more examples in Table 5); such trivial reasons should give rise to an intermediate label of ‘MAYBE’ and not ‘YES/NO’ - most of the cases where the annotators marked the rationales as unsatisfactory for property P3 were cases where Cave gave a ‘YES/NO’ label and the annotators deemed it should be a ‘MAYBE’.
Initial steps towards handling these issues include reduced reliance on linguistic features where Cave struggles such as punctuation style; this includes (1) making downstream users aware about them so that they can make an informed final decision with Cave’s rationales, (2) modifications to Cave’s prompt so that these features are used to make authorship conclusions only if the model observes unusual behaviour in that regard. As future work, we would like to address these shortcomings with targeted training. We would also like to look at dynamic weighing of linguistic features at inference time, i.e., which features contributes more/less to the authorship analysis of these two documents? Finally, we also discuss random errors owing to computational limitations (such as limited context window in LMs) in Appendix G.
We report these numbers for individual datasets in Table 12 in Appendix F. We also release the complete set of test-set rationales generated by Cave  [^9] Cave’s rationales Sheet.
4 Discussion
4.1 Accuracy vs. Explainability
Early works on self-rationalization such as Kojima et al. (2022); Marasovic et al. (2022) showed that rationale generation improved task accuracies; however, most such works place a higher focus on task performance than on the rationale quality. Recent works Joshi et al. (2023); Ramnath et al. (2023) discuss the need for increased rationale quality in order for them to be useful and trusted in downstream applications. Hence, in our efforts towards better Av, we place a high importance on rationale quality; Cave achieves both increased explainability (via consistency of labels) as well as competitive task accuracies.
We measured consistency. What next?
@@ 4 Discussion
In addition to consistency, there are several properties necessary for the rationale to be of good quality Joshi et al. (2023); Ramnath et al. (2023). Some of these properties are easily verifiable by humans, even if there exist no prevalent automatic methods to measure them: for eg:- factual correctness, hallucinated details in the rationale, etc. that we measured in our human evaluation (Section 3.5). There are also properties that are hard to verify even by human evaluators: for eg:- completeness of the rationale (i.e., has the rationale missed any significant similarity/dissimilarity between the documents?). In fact, completeness does not have a comprehensive definition even for simpler tasks like multi-choice QA Ramnath et al. (2023). In the future, we would like to work on automatic metrics and model training for these properties in Av.
@@ 4.2 Potential Risks
As discussed in Section 1, Av is a high-stakes task. While we strive to enforce explainability in Cave’s generations, it is still a language model which is trained on data generated from a black-box LM; both GPT-4-Turbo’s and LLaMa-3-8B’s biases could have been propagated to Cave. Further, as we see in Section 3.5, Cave’s rationales sometimes have hallucinated details, especially for certain linguistic features. Hence, we caution downstream users that Cave’s outputs should be used only after due understanding and verification. We discuss more details in the Limitations section at the end.
@@ 5 Conclusion
Existing methods for authorship verification face a lack of accessible explainability. In this work, we train an in-house language model Cave to generates structured and consistent explanations for Av. We present strong experimental results on three difficult Av datasets; we also present a human pilot study on the quality of generated rationales. Finally, we provide detailed discussions on explainability in Av and explore errors made by Cave due to factors such as hallucination, dataset biases, etc. to guide research in this area. As future work, we would like to explore reward metrics and reward-based learning for Av explanations.
@@ Acknowledgements
Like any large language model, Cave can unintentionally lead to toxic, incorrect or hallucinated text; it is the responsibility of the user to use the model mindfully. Further, since Cave is designed for the task of Av which is extremely sensitive in nature, it is the responsibility of the user to duly verify the data used to train the model, and also verify the output given by the model before using it. Further, as noted in the main paper, our models have not been explicitly trained for completeness of rationale; that is, apart from wrong information the generated explanation may have, it could also have missed some crucial similarity/dissimilarity between the documents.
Further, the Cave models trained in this paper were trained on data sampled from yet another language model GPT-4-Turbo and not on human-written gold standard data; it is possible that harmful biases present in GPT-4-Turbo (by process of its training and its training data) might have been passed over to the silver training data. It is necessary to keep this in mind while using our released training data.
Lastly we note that our method has practical limitations faced by any NLP task; our models are restricted by the size of the context window allowed by the model (both input and output text) (also keeping in mind restrictions due to computational costs, financial costs and GPU size available). This can lead to data missed by the model in its reasoning.
@@ 8. Other Aspects:
PromptAV (and by extension, CoT) which serve as the inspiration for Cave produce outputs which cannot be automatically checked for consistency; these prompts produces a free-text explanation followed by a final score which can be thresholded to produce a YES/NO Av label. As observed in the experiments of Hung et al. (2023), models tend to produce low confidence scores with PromptAV, which means that for better accuracy, the optimal threshold in their experiments with GPT-3.5-Turbo was found to be 0.2-0.3, which will most probably not be consistent with the free text rationale (that is, if the rationale says that it thinks it is not confident that the authors are the same, and the confidence is 0.25, the automatic label prediction will still assume that the predicted label is YES). Further, with just the score as the label, it is still possible that we misconstrue the label because the confidence score and the model’s predicted label in the text are different, as shown in the examples in Tables 8 and 9.
@@ Appendix C Hyperparameters
For our baseline SVM, we use TF-IDF feature vectors for Blog-Auth (words) and FanFiction (characters), and word-count features for IMDb62 (characters). All 3 datasets’ models use the RBF kernel. IMDb62 and FanFiction use n-grams of minimum size 1 and maximum size 2, whereas Blog-Auth uses only n-grams os size 2. IMDb62 uses a C-value of 5.0. We used sklearn version 1.2.2 for this baseline.
For our baseline of LLaMa-3-8B finetuned with CoT and PromptAV, our final training dataset sizes are: for IMDb62, Blog-Auth and FanFiction, the CoT sizes - 2.7k, 2.9k, 2.4k and the PromptAV sizes - 2.6k, 3k, 2.6k.
@@
Appendix D Cave with different offline models
Model Test $\rightarrow$ IMDb62 Blog-Auth FanFiction
Train $\downarrow$ Acc. Cons. Acc. Cons. Acc. Cons.
LLaMa-3-8B Single-dataset 75.7* 0.97 60.4 0.94 60.2 0.81
Multi-dataset 74.9 0.99* 57.9 0.96 58.6 0.91*
Mistral-7B Single-dataset 66.0 0.91 58.0 0.87 53.6 0.96
Multi-dataset 74.3 0.96 59.6 0.92 57.4 0.83
Stable-LM-Zephyr-3B Single-dataset 42.5 0.67 47.7 0.81 48.3 0.66
Multi-dataset 34.9 0.47 32.7 0.51 17.5 0.22
Gemma-2-2B Single-dataset 24.5 0.36 32.3 0.55 27.9 0.42
Multi-dataset 53.9 0.69 36.9 0.58 52.2 0.77
Table 13: Empirical Results comparing Cave trained with local models of different sizes. Mistral-7B is the next best performing model after LLaMa-3-8B. Gemma-2-2B and Stable-LM-Zephyr-3B however have lower performances; further, the former’s performance increases with multi-dataset training while the latter’s decreases.
We report experiments with local models of varying sizes in Table 13. After LLaMa-3-8B, Mistral-7B is the next best performing model. The smaller models Gemma-2-2B and Stable-LM-Zephyr-3B however have lower performances; further, the former’s performance increases with multi-dataset training while the latter’s decreases.
Appendix E Uncertainty of Cave
Our metric Consistency serves the dual purpose of measuring consistency as well as uncertainty of the model in its prediction. We mark that a rationale’s $\textsc{Consistency}=1$ if and only if the aggregated intermediate labels, final score and final label all match. If they don’t all match, we deem the output as inconsistent and hence uncertain. Hence, the consistency numbers in Table 3 reflect the uncertainty of the model in its output. Specifically, we note that Cave’s outputs show high consistencies (the highlighted/last two rows in Table 3) indicating that Cave is highly certain of its outputs.
Appendix F Human Evaluation